OpenStudy (anonymous):

A rocket travels 60 km vertically upward at a 90° angle from point A to point B. It then moves a distance of 60 km along the line segment BC, as shown below. ABCD is a trapezoid with AB equal to 60 kilometers and BC equal to 60 kilometers. A perpendicular to line CD is shown with a dashed line. This perpendicular passes through point B. The measure of the angle between BC and the perpendicular is 30 degrees. What is the height of the rocket, at point C, from the ground?

OpenStudy (anonymous):

I know that at some point I'll use an equation with the cotangent of 30, but other than that I am lost..

OpenStudy (anonymous):

The 30-60-90 ratio...the hypotenuse is 60. What's the shorter leg...?

OpenStudy (ganpat):

use trigonometry.. sin 30 = opp/ hyp SO opp = sin 30 * 60... = 30.. and from given, its rectangle so that length is 60.. so pint c = 60 + 30 = 90 Km

OpenStudy (anonymous):

The shorter leg would be 30.

OpenStudy (ganpat):

yup...

OpenStudy (anonymous):

Yes. Then the bottom figure is a rectangle, so 30 + 60 = ?

OpenStudy (anonymous):

Thanks, I feel stupid haha.

OpenStudy (anonymous):

np :)

OpenStudy (ganpat):

no worries, All the best !!!